OpenStudy (anonymous):

Solve 3x + 2y = –12 for y.

OpenStudy (anonymous):

\[2y=12-3x \rightarrow y=6-1,5 x\] You won't get a fixed result for y, but an infinity of results defined by that equation.

OpenStudy (saifoo.khan):

first get rid of x. 2y = -12 -3x now cross multiply. \[y = \frac{-12x-3x}{2} \]